# MDA: An Interpretable Multi-Modal Fusion with Missing Modalities and Intrinsic Noise

### v2

@@ Introduction
Derm7pt dataset (Kawahara et al. 2018) is a multi-modal dataset that contains clinical, dermoscopic, and meta-data modals for melanoma skin lesions in dermatology. We predict disease diagnosis (DIAG) and all 7 dermoscopic concepts, including blue-whitish veil (BWV), dots and globules (DaG), pigmentation (PIG), pigment network (PN), regression structures (RS), streaks (STR), and vascular structures (VS) in considered datasets. The dataset contains a total of 1011 cases.
@@
Multimodal IMDb (MM-IMDb) (Arevalo et al. 2017) contains two modalities: image and text. We conduct experiments on this dataset to predict a movie genre using image and text modality, which is a multi-label classification task as multiple genres could be assigned to a single movie. The dataset includes 25, 956 movies and 23 classes.
@@
Audiovision-MNIST (Vielzeuf et al. 2018) consists of an independent image and audio modalities. The images, which are digits from 0 to 9, are collected from the MNIST dataset (LeCun et al. 1998) with a size of 28 × 28, and the audio modality is collected from Free Spoken Digits Dataset containing raw 1500 audios. The dataset contains 1500 samples for both image and audio modalities.
Table 2: Multi-modal Fusion Results Under Perturbed Inputs with Ablation Experiments.
Type Modal Fusion Method Accuracy
Pre-trained Uni-modal EUS - 85.3
Uni-modal WLE - 81.7
Uni-modal Report - 87.4
Full Modalities Multi-modal EUS+WLE+Report w/o MDA 91.2
Multi-modal EUS+WLE+Report SCT Fusion (2023) 82.8
Multi-modal EUS+WLE+Report w/ MDA (Ours) 98.9
Missing Modalities Multi-modal Randomly missing 1 modal w/o MDA 88.1
Multi-modal Fixed missing 1 modal MMD (2022) 88.4
Multi-modal Randomly missing 1 modal w/ MDA (Ours) 95.4
Learning with Intrinsic Noise Uni-modal Report - 80.5
Multi-modal EUS+WLE+Report w/o MDA 89.3
Multi-modal EUS+WLE+Report w/ MDA (Ours) 97.9
    We simulated missing modalities by applying dropout to three different modalities. The full set of modalities was used in training, while during testing, only the modality with dropout was missing. The final result was obtained by averaging the predictions from the three models.
Experiments and discussion
In the first section, we present the results of our proposed method in addressing the challenges of modality heterogeneity, missing modalities, and learning with intrinsic noise in our gastrointestinal disease dataset. We comprehensively validate the effectiveness of our proposed method by analyzing three different fusion scenarios and the results of the ablation experiment. In the second section, We conducted experiments on publicly available datasets to evaluate the performance of the proposed method in multi-modal fusion and fusion efficacy in the presence of missing modalities. We compared our method with SOTA approaches for handling missing modalities, and in multiple tasks across multiple databases, MDA achieved SOTA results. The interpretability analysis of MDA is presented in the last section.
The efficacy of MDA in confronting the three key challenges of multi-modal fusion
Classifying Gastrointestinal tumors is crucial for surgical decision-making in clinical (Liu et al. 2021). Given their multi-modal characteristics, this paper will use submucosal tumor classification as an example to demonstrate our methodology. We evaluated the impact of EUS, WLE, and imaging reports in multi-modal fusion under three challenges, examining performance variations and the importance of each component in conducting ablation studies (see Table 2).
Modality Heterogeneity Challenge. The training of single-modal tasks is independent, with uni-modal features integrated into the MDA after model freezing. This is done to ensure that the self-attention module within the single-modal model does not engage in the learning process of the feature fusion module MDA, thereby eliminating any ambiguity in the analysis of MDA’s performance enhancement, as it does not affect the weight adjustments of MDA. The experimental results show that the direct concatenation-based fusion of multi-modal features achieved higher accuracy than any individual uni-modal testing accuracy, this indicates the effectiveness of multi-modal fusion in improving disease diagnosis accuracy on this dataset. By employing the MDA, which adaptively learns the weight relationships between modalities, the fusion capability of the multi-modal model was enhanced, resulting in a significant improvement in diagnostic accuracy compared to direct concatenation (acc: 91.2% $<$ 98.9%). This demonstrates the effectiveness of the proposed model in addressing modality heterogeneity.
The precision of MDA significantly surpasses that of other scalable multimodal methods such as SCT, demonstrating its advantage in handling mutual attention across modalities. MDA employs continuous attention learning without the need for additional branches, whereas SCT incurs a heavy computational burden and lacks sustained attention in fusion, resulting in suboptimal integration performance. Sustained attention is also a crucial factor that enables MDA to surpass the sustainable scalability in multimodality without augmenting the computational load of branching pathways.
Missing modality challenge. The results showed that the accuracy of direct concatenation fusion was lower than the accuracy of concatenation fusion using all modalities (acc: 88.1% $<$ 91.2%), indicating that the missing modality indeed affects the performance of multi-modal fusion, even with partial missing. However, when using MDA, even with only two modalities involved, the accuracy under attentional adjustment far exceeds that obtained with concat (accuracy: 95.4% $>$ 88.1%). This finding suggests that MDA can counteract most of the negative effects of missing modalities on feature fusion by shifting attention, as evidenced by the fact that it is only about 3% lower than the fusion accuracy using all modalities (acc: 95.4% $<$ 98.9%).
Learning with intrinsic noise challenge. The results showed that the wrong attribute significantly impacted the diagnostic accuracy of the model, as the accuracy decreased from 87.4% to 80.5%. After applying the concatenation-based multi-modal fusion, the accuracy improved to 89.3%. Although the image information helped mitigate the impact of erroneous modalities, the accuracy was still lower than that achieved with correct modalities (acc: 89.3% $<$ 91.2%). Upon incorporating the MDA, the model is capable of disregarding misleading information, as it establishes a correspondence between the report and image modalities. As a result, the diagnostic accuracy was essentially restored to the level achieved with the correct modalities (acc: 97.9%).
We compared the recent multi-modal fusion methods on a gastrointestinal dataset under the full modality setting. The results are shown in Table 2. For the SCT Fusion method (Hoffmann, Clasen, and Demir 2023), we computed the cross-entropy loss between our classification results and the ground truth of the eight-class labels without modifying the other configurations. The results demonstrate that the proposed method significantly outperforms recent multi-modal approaches in terms of fusion performance across multiple modalities (acc: 98.9% $>$ 82.8%). We observe that a pivotal reason for this outcome is the specialized nature of these methods, which have been designed for particular tasks, leading to reduced model portability. In contrast, the design flexibility of MDA enables its application across a broad spectrum of multimodal datasets.
@@
Table 3: Comparison of classification results to the state-of-the-art multi-modal fusion methods in terms of accuracy (%) by using Derm7pt (Kawahara et al. 2018) dataset.
@@
Baseline (Kawahara et al. 2018) 87.1 60.0 66.1 70.9 77.2 74.2 79.7 74.2 73.7
HcCNN (Bi et al. 2020) 87.1 65.6 68.6 70.6 80.8 71.6 84.8 69.9 74.9
FM4-SS (Tang et al. 2022) 88.1 66.1 70.1 71.1 81.5 78.0 81.8 78.5 77.0
RemixFormer (Xu et al. 2022a) - - - - - - - 81.3 -
CAFNet (He et al. 2023) 87.8 61.5 73.4 70.1 81.8 77.0 84.3 78.2 76.8
TFormer (Zhang, Xie, and Chen 2023) 86.7 66.9 70.3 74.3 82.1 76.7 83.0 79.5 77.4
MICA (Bie, Luo, and Chen 2024) 84.4 79.1 68.8 74.4 72.8 71.3 81.3 83.9 77.0
RGMC50 (Li et al. 2024) - - - - - - - 78.9 -
@@
Table 4: Model performance comparison of classification accuracy of missing modality (by setting different available audio rates) on avMNIST dataset. And train 30 epochs. ’M-m Concat’ refers to multi-modal fusion using concatenation, while ’Ours-m’ represents our proposed fusion method that handles missing modalities during testing. ’Ours-f’ denotes testing with full modalities, while only having nof the audio modality available in the training data.
Multi-Modal Rate Uni-Image Uni-Audio M-m Concat Ours-m SMIL (2021) ShaSpec (2023) Ours-f
70% 87.3 82.5 87.9 95.8 96.4 97.3 98.1
50% 87.3 82.5 84.3 86.3 95.6 96.0 96.2
20% 87.3 82.5 83.6 86.7 94.4 95.3 95.9
    Results for this case where only audio data is missing. Our approach involves random modalities being missing.
Comprehensive comparison in different database
In this chapter, we primarily utilize open datasets to conduct a comprehensive comparison of the proposed MDA method with the latest multi-modal fusion techniques across three major challenges. In the first section, we compared our MDA method with the latest techniques on the Derm7pt dataset, using a full-modal approach. In the second section, we compared MDA with the latest methods on two public multimodal datasets under the scenario of missing modalities.
@@
Comparison with related multi-modal fusion works To comprehensively evaluate MDA, we compared it with other state-of-the-art multi-modal fusion methods on the Derm7pt dataset, as shown in Table 3. The table reports the quantitative results of the classification accuracy for each concept, highlighting the performance of MDA on various tasks. MDA outperforms other methods in most metrics, particularly in disease diagnosis, where it achieves an accuracy of 85.7%, significantly surpassing the second-best method by 1.8%. This exceptional performance in disease diagnosis demonstrates the robustness and effectiveness of MDA in integrating and analyzing multi-modal data, which is crucial for accurate diagnosis in medical studies.
Table 5: Model performance comparison of classification accuracy of missing modality (by setting different available audio rates) on MM-IMDb dataset. Evaluating performance using F1 sample scores and training 5 epochs.
Multi-Modal Rate Uni-Image Uni-Text M-m Concat SMIL Ours-m
70% 0.361 0.398 0.637 - 0.701
50% 0.361 0.398 0.533 - 0.578
20% 0.361 0.398 0.496 0.541 0.546
    Results for this case where only text data is missing. Our approach involves random modalities being missing.
@@
Comparison with related missing modality works Due to the extensive research on missing modalities, we specifically conducted an extended investigation into the performance of MDA in the presence of missing modalities. Experiments were performed on the MM-IMDb and avMNIST datasets. We constructed datasets with 20%, 50%, and 70% of the full modality and randomized the missing modalities during training. We compared the accuracy of single-modal, direct multi-modal fusion, and the proposed method during evaluation and testing. We also tested the proposed model’s proficiency in multi-modal fusion using 100% images and naudio during training, comparing the results with SMIL (Ma et al. 2021) and ShaSpec (Wang et al. 2023).
Table 4 and Table 5 present experimental results on avMNIST and MM-IMDb datasets. The results indicate that when only 20% of the multimodal data is included, MDA may underperform compared to single-modality results due to the lack of sufficient information to learn the linear attention relationships between the two modalities. However, the results are still superior to concatenation-based fusion. When evaluating the influence of varying multi-modal rates on accuracy, MDA exhibits a markedly superior resistance to interference compared to the concatenation fusion method. This suggests that MDA is proficient in leveraging limited information, which is a design strength of MDA, enabling the adaptive allocation of weights to information from different modalities. Table 4 presents the performance of Ours-f when tested with both audio and image modalities, with only nof the audio available in training. MDA outperforms SMIL and ShaSpec, especially when 70% of the audio is missing. This is because MDA’s ability to adaptively transfer attention from missing modalities to others preserves feature reliability, whereas other methods may compromise feature quality through generated missing modalities. Moreover, MDA’s versatility in handling any missing modality makes it more applicable to real-world scenarios.
Figure 3: Macroscopic investigation of attention weights for various diseases
Interpretability analysis
In medical research, attention visualization methods like CAM are used for interpretability but have limitations in analyzing model attention across modalities and disease categories. To address this, we use adaptive attention maps across modalities to comprehensively analyze model performance in multi-modal fusion, modality absence, and handling intrinsic noise. Interpretability in Diagnostic Standards Research In clinical practice, multi-modal imaging techniques are frequently employed to gather comprehensive diagnostic information. For instance, the combination of EUS, WLE, and CT/MRI is utilized for the diagnosis of gastrointestinal diseases, with each imaging modality being sensitive to different features. Different diseases necessitate attention to various imaging modalities, with some having established gold standards for diagnosis, such as WLE for Lipoma. However, many diseases lack a definitive gold standard  (Richardson et al. 1995; Chandler et al. 2019; Jacobson et al. 2023). The increasing medical focus on establishing diagnostic gold standards aims to standardize procedures and decrease required imaging modalities. The study seeks to provide evidence for definitive gold standards and serves as a reference for diseases without them. It is the pioneering research in computer-aided diagnosis to investigate the interpretability of detailed features and gold standard modalities, with the goal of enhancing AI’s role in setting diagnostic standards.

### v3

@@ 3 Datasets
Gastrointestinal disease dataset We compiled a multi-center gastrointestinal disease dataset comprising three modalities: Endoscopic Ultrasound (EUS), White Light Endoscopy (WLE), and imaging reports. The dataset was gathered from two separate centers (***hospital and ***hospital) and captured using Olympus UM-2R and UM-3R equipment, totaling 421 cases spanning eight distinct categories (Gastrointestinal Stromal Tumor (GIST), Gastric Leiomyoma (GLM), Neuroendocrine Tumor (NET), Ectopic Pancrea (EP), Lipoma, Gastrointestinal Schwannomas (GS), Pneumatosis Cystoid (PC), and Inflammatory Fibroid Polyp (IFP)). Among these patients, the gender distribution is 53% male to 47% female, with an age range spanning from 27 to 78 years. The imaging reports (in Chinese) offer detailed descriptions of lesion characteristics observed within the two imaging modalities. The detailed data quantities for each category are shown in Table 1. We divided the dataset into training and test sets based on the variability of the centers, encompassing 327 and 94 cases, respectively.
@@
Derm7pt dataset [19] is a multi-modal dataset that contains clinical, dermoscopic, and meta-data models for melanoma skin lesions in dermatology. We predict disease diagnosis (DIAG) and all 7 dermoscopic concepts, including blue-whitish veil (BWV), dots and globules (DaG), pigmentation (PIG), pigment network (PN), regression structures (RS), streaks (STR), and vascular structures (VS) in considered datasets. The dataset contains a total of 1011 cases.
@@
Multi-modal IMDb (MM-IMDb) [1] contains two modalities: image and text. We conduct experiments on this dataset to predict a movie genre using image and text modality, which is a multi-label classification task as multiple genres could be assigned to a single movie. The dataset includes 25, 956 movies and 23 classes. The division rule for this dataset adheres to the contrasting method SMIL [28].
@@
Audiovision-MNIST [41] consists of an independent image and audio modalities. The images, which are digits from 0 to 9, are collected from the MNIST dataset [21] with a size of 28 × 28, and the audio modality is collected from Free Spoken Digits Dataset containing raw 1500 audios. The dataset contains 1500 samples for both image and audio modalities. The division rule for this dataset adheres to the contrasting method SMIL [28].
4 Experiments and discussion
In our experiments, each uni-modal task is trained independently, with its parameters frozen during multi-modal fusion training. This approach ensures that the self-attention module of each uni-modal model does not interfere with the fusion module’s learning process, allowing any accuracy improvements to be attributed solely to the different multi-modal fusion strategies.
Table 2: Ablation experiments of multi-modal fusion assessing diagnostic accuracy (%) under conditions of missing modalities and intrinsic noise.
Fusion Intrinsic Noise
Modal Method w/o 5% 20%
Uni-EUS - 85.3 - -
Uni-WLE - 81.7 - -
Uni-Report - 87.4 85.7 80.2
EUS+WLE MDA (Ours) 91.6 - -
EUS+Report MDA (Ours) 95.2 - -
WLE+Report MDA (Ours) 94.7 - -
Complete Modalities Addition 87.6 87.0 85.5
Maximum 88.3 88.2 86.7
Concatenate 91.2 89.3 87.8
MDA (Ours) 98.9 98.6 98.5
70% Modalities Addition 87.3 85.6 84.1
Maximum 87.4 86.3 85.6
Concatenate 89.7 88.4 87.9
MDA (Ours) 98.6 98.4 97.3
50% Modalities Addition 85.7 82.8 81.9
Maximum 84.2 81.1 80.2
Concatenate 89.1 86.3 85.2
MDA (Ours) 97.4 97.1 96.5
20% Modalities Addition 82.3 80.2 79.6
Maximum 81.9 80.4 79.3
Concatenate 84.7 82.1 81.0
MDA (Ours) 95.6 95.7 94.6
@@ 4 Experiments and discussion
Table 3: Comparison of classification results to the state-of-the-art multi-modal fusion methods in terms of accuracy (%) by using Derm7pt [19] dataset. The best values are highlighted in bold, and the second-best ones are highlighted underlined.
@@
Baseline [19] 87.1 60.0 66.1 70.9 77.2 74.2 79.7 74.2 73.7
HcCNN [3] 87.1 65.6 68.6 70.6 80.8 71.6 84.8 69.9 74.9
FM4-SS [39] 88.1 66.1 70.1 71.1 81.5 78.0 81.8 78.5 77.0
RemixFormer [47] - - - - - - - 81.3 -
CAFNet [11] 87.8 61.5 73.4 70.1 81.8 77.0 84.3 78.2 76.8
TFormer [51] 86.7 66.9 70.3 74.3 82.1 76.7 83.0 79.5 77.4
MICA [4] 84.4 79.1 68.8 74.4 72.8 71.3 81.3 83.9 77.0
RGMC50 [23] - - - - - - - 78.9 -
@@
4.1 The efficacy of MDA in confronting the three key challenges of multi-modal fusion
We present the results of our proposed method in addressing the challenges of multi-modal fusion, random missing modality, and learning with intrinsic noise in our gastrointestinal disease dataset. Table 2 shows the efficacy of our method for the three challenges via extensive analysis of ablation test results across three fusion scenarios.
Scalability for multi-modality integration challenge. The experimental results underscore the importance of multi-modal integration in enhancing diagnostic accuracy. Initially, when considering uni-modal approaches, the performance is limited, with individual modalities such as Uni-EUS and Uni-Report achieving accuracies of 85.3% and 87.4%, respectively. This highlights the inherent limitations of relying on a single data stream for diagnosis. As we progress to multi-modal fusion, the accuracy significantly improves, indicating the complementary nature of different modalities. For instance, the simple Concatenate fusion method already shows an enhancement, with an accuracy of 91.2% when all modalities are present, surpassing the accuracy of any single modality.
As the number of integrated modalities increases, the MDA method exhibits a consistent and substantial improvement in accuracy. With the addition of each modality, the MDA’s accuracy climbs from 95.2% with two modalities to 98.9% with three modalities, demonstrating its ability to effectively integrate and leverage information from multiple sources without increasing the computational load associated with branching pathways. This sustained attention mechanism in MDA is pivotal for achieving superior sustainable scalability in multi-modal data processing. This robustness in scalability, combined with its efficient approach to attention, allows MDA to excel in precision and scalability, making it a more effective method for integrating and processing multi-modal data in real-world scenarios.
Missing modality challenge. The sensitivity of multi-modal fusion methods to missing modalities is a critical factor in their practical application. Traditional fusion techniques such as concatenation or Addition often struggle to maintain performance when faced with incomplete data. In contrast, MDA exhibits remarkable stability in the face of missing modalities. The impact of missing data on MDA’s performance is substantially mitigated by its advanced attentional mechanism. When 70% of the modalities are present, MDA still achieves an impressive accuracy of 98.6%, which is a mere 0.3% decrease from its performance with all modalities. This minimal change in accuracy, even with a large proportion of modalities missing. Further examining the resilience of MDA, we observe that as the percentage of missing modalities increases, the performance of other fusion methods tends to degrade more significantly. For example, if 80% of one modality is randomly missing, the concatenate fusion accuracy might drop to 84.7%, while MDA could still maintain an accuracy close to 95.6%. This stark contrast in performance underscores the superior stability of MDA when dealing with substantial data gaps.
The stability of MDA in the face of missing modalities is not just a matter of maintaining high accuracy; it also reflects the method’s ability to effectively weigh and integrate the available information. The attention mechanism in MDA is designed to dynamically adjust to the presence or absence of modalities, ensuring that the model can leverage the most relevant and informative parts of the data. This adaptability is crucial in real-world diagnostic settings where the completeness of data cannot be guaranteed, and MDA’s performance under these conditions makes it a standout choice for multi-modal data integration.
Learning with intrinsic noise challenge. Table 2 highlights the resilience of MDA to intrinsic noise. While the accuracy of the Uni-Report modality decreases from 87.4% without intrinsic noise to 80.2% with 20% noise, MDA maintains an impressive accuracy of 98.5% with the same level of noise. This is a stark contrast to the concatenate method, which sees a decrease from 91.2% to 87.8% with 20% noise. The MDA method’s ability to establish correspondences between modalities and effectively diminish the contribution of noisy modalities to the decision-making process is evident, as it restores diagnostic accuracy to a level comparable to that achieved with noise-free data (98.5% vs. 98.9%).
The analysis of the interplay between missing modalities and inherent noise has illuminated the resilience of different fusion techniques in the face of intricate modal forms. The results show that as the proportion of missing modalities increases, all methods experience a decline, but MDA consistently outperforms the others, even in noisy conditions. For example, with 50% missing modalities and 5% noise, MDA achieves 97.1%, while traditional methods drop to 82.8%, 81.1%, and 86.3%. Even with 20% noise, MDA’s accuracy only slightly decreases to 96.5%, while others see sharper declines. MDA’s superior performance is due to its advanced continuous attention algorithms, which allow it to handle incomplete and noisy data more effectively than traditional methods. As noise levels rise, the performance gap widens, particularly with 50% and 20 modality loss. The sharp decline in the efficacy of traditional methods highlights their vulnerability to noise when faced with substantial gaps in the data, whereas MDA demonstrates a more stable and resilient performance, underscoring its superior ability to maintain diagnostic accuracy even in the presence of substantial noise and missing data. This pronounced discrepancy underscores the robustness of MDA in handling complex data challenges, reaffirming its advantage over traditional approaches in the realm of multi-modal data integration.
4.2 Comprehensive comparison in different database
We conducted experiments on publicly available datasets to assess the efficacy of our proposed method in both complete and missing modal fusion scenarios. We benchmarked our approach against SOTA methods tailored for handling missing modalities. Across a range of tasks and various databases, MDA demonstrated SOTA performance.
@@ 4.2 Comprehensive comparison in different database
Comparison with related multi-modal fusion works The comparison results on the Derm7pt dataset are shown in Table 3. The table reports the quantitative results of the classification accuracy for each concept, highlighting the performance of MDA on various tasks. MDA outperforms other methods in most metrics, particularly in disease diagnosis, where it achieves an accuracy of 85.7%, significantly surpassing the second-best method by 1.8%. This exceptional performance in disease diagnosis demonstrates the robustness and effectiveness of MDA in integrating and analyzing multi-modal data, which is crucial for accurate diagnosis in medical studies.
@@
Comparison with related missing modality works Due to the extensive research on missing modalities, we specifically conducted an extended investigation into the performance of MDA in the presence of missing modalities. Experiments were performed on the MM-IMDb and avMNIST datasets.
Table 4: Model performance comparison of classification accuracy of missing modality on avMNIST dataset. $\eta\%$ represents the rates of complete modalities included.
Method 100% 70% 50% 20%
Uni-Image 87.3 - - -
Uni-Audio 82.5 - - -
AutoEncoder [2] - - - 88.8
GAN [10] - - - 89.5
Full2miss [37] - - - 92.6
SMIL [28] 97.6 96.4 95.6 94.4
ShaSpec [43] 98.2 97.3 96.0 94.6
MCKD [44] - - - 95.1
Concatenate 92.7 88.1 84.7 84.2
MDA (Ours) 98.4 98.1 96.2 95.9
  • Only audio modality is missing.
  • Missing modality at random.
Table 5: Model performance comparison of classification accuracy of missing modality on MM-IMDb dataset. Evaluating performance using F1 sample and F1 Macro scores. $\eta\%$ represents the rates of complete modalities included.
F1 sample F1 Macro
100% 70% 20% 30% 30%
Method w/o Img w/o Text
Uni-Image 36.1 - - - -
Uni-Text 39.8 - - - -
MFAS [33] 63.1 - - - -
CentralNet [41] 63.9 - - - -
ViLT [20] 64.6 - - - -
SMIL [28] - - 54.1 - -
MAPs [22] - - - 46.3 39.2
MuAP [7] - - - 47.2 41.3
MSPs [15] - - - 47.4 38.3
Concatenate 66.2 63.7 49.6 40.4
MDA (Ours) 71.8 70.1 54.6 47.7
    Only text modality is missing.
  • $\dagger$
    Fixed missing modality. Different prompts are set for different missing modalities.
  • $\ddagger$
    Missing modality at random.
The experimental results for both the avMNIST and MM-IMDb datasets, as detailed in Tables 4 and 5, underscore the robustness and adaptability of MDA. In the avMNIST dataset, MDA excels across all levels of data availability, from 20% to 100%, outperforming traditional methods like Concatenate and AutoEncoder, as well as state-of-the-art techniques such as SMIL and ShaSpec. This success is attributed to MDA’s ability to dynamically allocate weights to available modalities, ensuring high classification accuracy even with limited data. A similar trend is observed in the MM-IMDb dataset, where MDA achieves the highest F1 sample and F1 Macro scores, maintaining superior performance with missing modality, particularly when 70% of modalities are randomly included. Notably, MDA also demonstrates a clear advantage over methods like MFAS, CentralNet, and ViLT which depend on complete modalities. In the MM-IMDb dataset, the performance gap between MDA and methods tailored for fixed missing modalities, such as MAPs, MuAP, and MSPs, under random missing conditions, further emphasizes MDA’s adaptability and generalizability. Overall, MDA’s consistent and superior performance across varying conditions of data completeness highlights its potential for real-world applications, particularly in medical diagnostics where data integrity is often variable.
4.3 Interpretability analysis
To address Challenge 4: The Comprehensive Interpretability of Multi-modal Fusion, we have conducted a thorough analysis of the interpretability of multi-modal fusion on MDA from both a macroscopic clinical diagnostic prior perspective and a microscopic model perspective, focusing on the variations in multi-modal attention.
Figure 4: Macroscopic investigation of MDA weights for various diseases.
Interpretability in Diagnostic Standards Research. Fig. 4 illustrates the average MDA weights of the proposed method for different disease categories. MDA weights $w_{i}$ for the modality $i$ can be calculated as in Eq. 8, where ${\mathbf{\Theta}_{i}}$ represents the MDA map, as discussed in Section 2.4.
$w_{i}=\frac{\sum\mathbf{\Theta}_{i}}{\sum_{i=1}^{N}\sum\mathbf{\Theta}_{i}}$ (8)